# Wonder3D++: Cross-domain Diffusion for High-fidelity 3D Generation from a Single Image

### v1

@@ Appendix B Effectiveness of cross-domain multi-view enhancement module
Figure 14: Quantitative comparison with Direct2.5 [44]
@@ Appendix C More compassion on single image 3D mesh generation
Figure 17: Qualitative comparison on single image to 3D task between InstantMesh [82] The mesh generated by our method not only has better geometric and texture details but also has higher consistency with the input image.
In this section, we showcase more results on single image 3D mesh generation task compared with InstantMesh [82] in figure 17. Our method generates meshes with better geometric details and texture fidelity.
@@ Appendix D Cross-domain novel view synthesis
Figure 19: Qualitative comparison on cross-domain NVS task with Wonder3D [41]. Our method generates more fidelity rgb images and sharper normal maps.
@@ Appendix G Text-to-3D Generation
While primarily designed for image-to-3D reconstruction, Wonder3D++’s architecture allows for seamless integration with text-to-image diffusion models, such as Stable Diffusion [56], to enable text-driven 3D generation. This is achieved by first generating a single image conditioned on the input text and subsequently processing it through our pipeline to produce a 3D mesh. As demonstrated in Fig. 14, our method generates 3D assets exhibiting superior texture fidelity and geometric accuracy when compared to Direct2.5 [44]. In contrast to our cross-domain joint sampling scheme, Direct2.5 [44] employs a two-stage pipeline involving multi-view normal map generation followed by image generation for mesh extraction. This separation hinders the model from learning a coherent texture and geometry data distribution. A comprehensive numerical comparison is presented in Table V.
@@
To evaluate the fidelity, detailed geometry, and texture quality of 3D meshes generated by our method, as well as its generalization capabilities, we conduct visual comparisons on examples from the GSO [14] dataset against both ground truth (GT) meshes and results from InstantMesh [82]. As depicted in Fig. 16, the meshes generated by our method maintain a high degree of similarity to the ground truth, while exhibiting only minor loss in fine details.

### v2

@@ Appendix B Effectiveness of cross-domain multi-view enhancement module
Figure 14: Quantitative comparison with Direct2.5 [lu2024direct2]
@@ Appendix C More compassion on single image 3D mesh generation
Figure 17: Qualitative comparison on single image to 3D task between InstantMesh [xu2024instantmesh] The mesh generated by our method not only has better geometric and texture details but also has higher consistency with the input image.
In this section, we showcase more results on single image 3D mesh generation task compared with InstantMesh [xu2024instantmesh] in figure 17. Our method generates meshes with better geometric details and texture fidelity.
@@ Appendix D Cross-domain novel view synthesis
Figure 19: Qualitative comparison on cross-domain NVS task with Wonder3D [long2024wonder3d]. Our method generates more fidelity rgb images and sharper normal maps.
@@ Appendix G Text-to-3D Generation
While primarily designed for image-to-3D reconstruction, Wonder3D++’s architecture allows for seamless integration with text-to-image diffusion models, such as Stable Diffusion [rombach2022high], to enable text-driven 3D generation. This is achieved by first generating a single image conditioned on the input text and subsequently processing it through our pipeline to produce a 3D mesh. As demonstrated in Fig. 14, our method generates 3D assets exhibiting superior texture fidelity and geometric accuracy when compared to Direct2.5 [lu2024direct2]. In contrast to our cross-domain joint sampling scheme, Direct2.5 [lu2024direct2] employs a two-stage pipeline involving multi-view normal map generation followed by image generation for mesh extraction. This separation hinders the model from learning a coherent texture and geometry data distribution. A comprehensive numerical comparison is presented in Table V.
@@
To evaluate the fidelity, detailed geometry, and texture quality of 3D meshes generated by our method, as well as its generalization capabilities, we conduct visual comparisons on examples from the GSO [downs2022google] dataset against both ground truth (GT) meshes and results from InstantMesh [xu2024instantmesh]. As depicted in Fig. 16, the meshes generated by our method maintain a high degree of similarity to the ground truth, while exhibiting only minor loss in fine details.